OpenStudy (some.random.cool.kid):

What is the equation in point−slope form of the line passing through (−3, −4) and (0, 2)? (y − 8) = 2(x − 3) (y + 3) = −2(x + 4) (y − 2) = −2(x + 4) (y + 4) = 2(x + 3)

OpenStudy (some.random.cool.kid):

@AloneS @zepdrix

OpenStudy (some.random.cool.kid):

@JoannaBlackwelder @Jaynator495

OpenStudy (amit1235813):

Step 1: Start from the standard equation of a straight line

OpenStudy (sshayer):

\[y-y1=\frac{ y2-y1 }{ x2-x1 }(x2-x1)\]

OpenStudy (some.random.cool.kid):

?? @sshayer

OpenStudy (some.random.cool.kid):

can you simplify that?

OpenStudy (some.random.cool.kid):

denominator?

OpenStudy (some.random.cool.kid):

2

OpenStudy (sshayer):

\[y-(-4)=\frac{ 2-(-4) }{ 0-(-3) }(x-(-3))\]

OpenStudy (some.random.cool.kid):

ok I see.. Im going to say (A)... once it factors out.

OpenStudy (some.random.cool.kid):

correct?

OpenStudy (sshayer):

\[y+4=\frac{ 6 }{ 3 }(x+3)\]

OpenStudy (sshayer):

can you simplify it?

OpenStudy (some.random.cool.kid):

y + 4 = 2(x+3).. divide whats on the out of the parin?

OpenStudy (some.random.cool.kid):

i think i got it thanks